### v1

# Do Construction Distributions Shape Formal Language Learning In German BabyLMs?

Abstract
We analyze the influence of utterance-level construction distributions in German child-directed speech on the resulting formal linguistic competence and the underlying learning trajectories for small language models trained on a novel collection of developmentally plausible language data for German. We find that trajectories are surprisingly robust for markedly different distributions of constructions in the training data, which have little effect on final accuracies and almost no effect on global learning trajectories. While syntax learning benefits from more complex utterances, lexical learning culminates in better scores with more fragmentary data. We argue that LMs trained on developmentally plausible data can contribute to debates on how rich or impoverished linguistic stimuli actually are.
Do Construction Distributions Shape Formal Language Learning In German BabyLMs?
Bastian Bunzeck  and Daniel Duran  and Sina Zarrieß CRC 1646 – Linguistic Creativity in Communication Department of Linguistics Bielefeld University, Germany {firstname.lastname}@uni-bielefeld.de
1 Introduction
One of the most contentious issues in language acquisition is the relationship between the input that the learner receives and the resulting linguistic system. Child-directed speech (or CDS) is structurally simple and, especially in the first three years of life, abounds with questions, imperatives, and fragmentary utterances, but features very few complex sentences Cameron-Faulkner et al. (2003); Bunzeck and Diessel (2024). This distribution of utterance-level constructions is conducive to the functional side of language acquisition: caregivers talk in this way to elicit responses, steer behavior or establish joined attention. But how do children acquire full-fledged, formal grammatical knowledge from such supposedly skewed input? Generativist approaches see this kind of stimulus as too impoverished to kickstart formal linguistic development and assume an underlying, innate grammatical endowment that enables it (cf. Chomsky, 1965; Crain and Pietroski, 2001; Thomas, 2002; Berwick et al., 2011), whereas usage-based scholars argue that domain-general learning mechanisms are sufficient to develop a full-fledged mental linguistic system from this input, which is rich enough (cf. MacWhinney, 2004; Tomasello, 2005).
The connectionist “renaissance”, fueled by the increased adoption of deep learning and transformer language models, has opened up new avenues of addressing such questions Warstadt and Bowman (2022). Developmental AI Dupoux (2018) and the BabyLM community Warstadt et al. (2023); Hu et al. (2024); Charpentier et al. (2025) have demonstrated that supposedly tabula rasa learners can indeed acquire syntactic structures, the formal side of language, successfully from very little data Huebner et al. (2021), but without tapping into the functional side of language (Mahowald et al., 2024). This makes them ideal testbeds for the aforementioned issue: is the construction distribution found in child-directed speech really too impoverished to learn formal linguistic capabilities, and are other, more complex registers, e.g. book text Cameron-Faulkner and Noble (2013); Noble et al. (2018), a richer and more beneficial form of input?
To investigate this, we compile a novel German BabyLM training set, for which we conduct the first utterance-level construction analysis for German. We find that distributions align with findings for English and other langauges, and that child-directed media (e.g., children’s encyclopedias) provide a construction-level middle ground between CDS and more complex book text. We then create three 5M-token subsets with distinct constructional profiles, e.g., varying the proportion of fragmentary and complex utterances, and train small, character- and subword-level Llama models on them. Our evaluation with lexical, syntactic, and semantic minimal pairs Bunzeck et al. (2025); Mueller et al. (2020); He et al. (2025) reveals that differences between grammatically complex training data and a developmentally plausible constructional distribution are fairly small. While certain syntactic phenomena are learned somewhat better from more complex sentences, lexical learning improves with more fragments and questions in the input. Most interestingly, input complexity only modulates the steepness of the resulting learning trajectories, but has no principal effect on the amount of input needed to kickstart learning.
2 Constructions in children’s input
Child-directed speech can be seen as a separate linguistic register and is the primary input that children encounter in their first years. It has mostly been scrutinized from phonetic and lexical viewpoints, where well-established findings are that it features slower speech and exaggerated intonation patterns which infants prefer listening to Zangl and Mills (2007), while its vocabulary is rather restricted to everyday topics and children’s immediate surroundings Snow and Ferguson (1977). Structurally, child-directed utterances are usually shorter and simpler than adult-directed ones Genovese et al. (2020) and feature high amounts of structural and lexical repetition Tal et al. (2024). Statistical properties of the input directly influence the children’s order of acquisition for syntactic patterns Huttenlocher et al. (2002); Ambridge et al. (2015), e.g. for relative clauses Diessel and Tomasello (2000); Brandt et al. (2008); Chen and Shirai (2015).
Early studies were mostly concerned with mapping out how much CDS is ungrammatical or otherwise “wrong” (in the sense of hesitations, false starts, etc., cf. Pine, 1994), but the quantitative turn in linguistics Janda (2013) has enabled more holistic analyses. In a seminal study, Cameron-Faulkner et al. (2003) analyze utterance-level constructions in child-directed English via a large corpus of toyplay sessions featuring children and their caregivers. They show that CDS features only few “canonical” SV(X)-utterances but abounds with questions, lexical fragments, or copula constructions. The reported construction distributions also hold for typologically different languages, e.g., Irish Cameron-Faulkner and Hickey (2011). These constructions and their real-world functions help children to quickly understand the functional side of language. On the other hand, the most common and repetitive utterances that English-speaking children hear from their mothers represent a rather skewed sample of the presumed, underlying formal language system. Generativist approaches would argue that certain formal structures, e.g. question formation from relative clauses, are not attainable from this kind of language, as the input never contains specific examples Chomsky (1980) (although Pullum and Scholz (2002) argue that the input frequently contains exactly such specific examples). Conversely, constructivist approaches, which view language learning as the re-construction of the target language Behrens (2021), argue that this kind of input is actually conducive to formal aspects of acquisition, by providing anchor points for first words and their semantic links to real-world reference, which then serve as building blocks for a gradual development into larger schemas (like questions with relative clauses).
Although CDS features such a skewed construction distribution, written language aimed at children, i.e., in the form of children’s books, is characterized by a much higher rate of canonical SV(X)-constructions than found in CDS Cameron-Faulkner and Noble (2013). In contrast, questions rarely occur in books. CDS produced in shared book reading presents a middle-ground — it contains more complex and SV(X)-constructions than regular CDS, but less than book text alone Noble et al. (2018). They argue that shared reading therefore plays an important role in moving children from early, isolated traces of linguistic knowledge to a rich mental language system. This also aligns with the findings by Bunzeck and Diessel (2024), who show that the distribution of constructions in CDS varies with situation type (toyplay features most questions, meal sessions beget more imperatives, shared book reading features more complex constructions) and child age (questions and imperatives become less frequent with age). They suggest that CDS is therefore adapted to support children’s cognitive and linguistic development. Yet, as corpus studies are necessarily descriptive and cannot establish causal/mechanistic connections on their own (e.g. what would happen if a child never hears CDS), it remains questionable if this is actually true.
3 Input in developmentally plausible LMs
Authentic data
Early approaches to modeling language acquisition with neural networks used hand-picked, manually ordered data points Rumelhart and McClelland (1986) or synthetic data generated with hand-crafted grammars Elman (1993); Christiansen and Chater (1999); Chang et al. (2006). Both lack developmental plausibility. Since then, data availability has improved with the establishment of developmental corpora. Frequently, CDS from the CHILDES database MacWhinney (2000) is used to train developmentally plausible LMs (cf. Pannitto and Herbelot, 2020; Huebner et al., 2021). While CHILDES-based models have the advantage of learning from authentic data only, they have the disadvantage of not accessing the full breadth of the linguistic input children receive. Children are exposed to many more different registers of language throughout their linguistic development, like the aforementioned shared (or solitary) book reading, or television shows Montag (2019); Gowenlock et al. (2024). As a response to this, the BabyLM corpora propose a data mix of varied spoken and written sources, from CDS over adult-adult conversations to OpenSubtitles Lison and Tiedemann (2016), but also children’s Hill et al. (2015) and adults’ books Gerlach and Font-Clos (2020). All data included in them could be plausibly encountered by children, which provides opportunities to ablate the influence of architecture/training on the learned linguistic knowledge.
For languages other than English, data availability is the greatest problem for the composition of developmentally plausible datasets. Salhan et al. (2024) use only data available from CHILDES for French, German, Japanese, and Chinese models, whereas Prévot et al. (2024) compare models trained on spoken data (child-directed + adult-adult conversations) with models trained on the French Wikipedia. As such, these first forays into more polyglot BabyLMs are still constrained to the child-directed input found in CHILDES and do not extend to the variety of inputs that children are exposed to Soderstrom (2007); Gowenlock et al. (2024). Notably, Suozzi et al. (2025) introduce an Italian BabyLM but do not elaborate on their data sources beyond CHILDES.
Linguistic properties
The actual linguistic make-up of pre-training corpora and its influence on linguistic performance have only recently begun to receive increased scrutiny. Focusing on the lexical level, Yam and Paek (2024) measure sentence-level textual complexity with readability metrics based on text-wide word/syllable–sentence ratios for different corpora (CHILDES only, BabyLM corpus, synthetic data, Project Gutenberg). They find that models trained on more complex text perform better at syntactic benchmarks, but simpler data (CHILDES) is learned better in terms of perplexity and loss convergence. Also on the lexical level, Muckatira et al. (2024) filter regular, non-BabyLM pre-training corpora for text spans that only contain vocabulary also found in English CHILDES data and find that simplified models generate more coherent text than models trained on more complex data and also succeed in syntactic tests if the test data is filtered accordingly. In contrast, Edman et al. (2024) change the semantic content of the pre-training data and use datasets that approximate linguistic input that second-language learners get, e.g., dictionary entries, grammar books, and paraphrased sentences. While grammar books moderately improve syntactic evaluation, there is no positive effect observable for the addition of the other text types.
Dataset Description # Words
CHILDES MacWhinney (2000) Child-directed speech 3,626,301
Child speech 1,511,144
OpenSubtitles Lison and Tiedemann (2016) Movie subtitles 1,543,094
CallHome Karins et al. (1997) Phone conversations 176,313
Klexikon Children’s online encyclopedia 1,384,891
MiniKlexikon Simplified online encyclopedia 272,886
Wikibooks Wikijunior Educational books 226,773
Fluter German youth magazine 2,862,278
Project Gutenberg Literature (children’s and young adult) 2,476,133
Dreambank Domhoff and Schneider (2008) Dream reports 939,197
Leipzig corpus news texts Goldhahn et al. (2012) Short news texts 1,541,803
Total 16,560,813
Table 1: Lexical token counts for all subcorpora of our corpus
Filtered corpora
While actual research on the structural/syntactic properties of the input is still rather sparse, training on filtered corpora has been used in pilot studies. Patil et al. (2024) and Misra and Mahowald (2024) filter out specific grammatical constructions from the standard BabyLM corpora and then probe the resulting models for knowledge of these grammatical constructions (which might also be analogically learned from related constructions or constructed from their parts). Patil et al. (2024) show that their models succeed on the BLiMP benchmark Warstadt et al. (2020), even if sentences containing structures targeted in BLiMP’s minimal pair sets are removed. Similarly, Misra and Mahowald (2024) show that acceptability scores for the English AANN construction can be reliably estimated from models that have never seen this specific construction. In sum, then, models appear to be able to generalize from indirect evidence and learn language in a somewhat constructivist, bottom-up fashion.
So far, previous work in NLP has failed to adequately investigate the structural composition of child-directed data. Most studies focus on lexical or semantic properties, emphasizing content over structure, and child-directed data is usually equated with a somewhat fitting vocabulary or with just being authentic data. However, findings from usage-based linguistics suggest that structural properties, like utterance-level construction distributions, play a crucial role in language acquisition. Patil et al. (2024) and Misra and Mahowald (2024) remove specific constructions from their data, whereas we aim to explore whether different global distributions of constructions influence the resulting linguistic knowledge and learning trajectories.
4 A German BabyLM dataset
Figure 1: Proportions of utterance-level constructions for all subcorpora in our corpus
To construct a German dataset, we use a variety of developmentally plausible sources, similar to the English BabyLM data Warstadt et al. (2023); Choshen et al. (2024). We use i) all data from German CHILDES corpora MacWhinney (2000), including frog stories from TalkBank Berman and Slobin (1994) and math lessons from ClassBank Stigler et al. (2000), ii) subtitles from OpenSubtitles Lison and Tiedemann (2016), iii) adult conversations from the CallHome corpus Karins et al. (1997), and iv) written data from Project Gutenberg, from which we downloaded a manually curated sample of children’s books, young adult literature and literature commonly read in German schools. We supply this data with two corpora, the DreamBank database of self-reported dreams Domhoff and Schneider (2008) and short news texts from the Leipzig corpus Goldhahn et al. (2012), although they are not child-directed per se, these sources are child-available in everyday language.
To approximate child-available input even better, we tap into freely available child/learner-directed sources and compile four additional subcorpora for our dataset. The Wikibooks Wikijunior shelve features educational resources aimed at children, focusing on a diverse array of topics such as technology or nature. The Klexikon is a children’s wiki in German, featuring more than 3,000 articles aimed at children between 5–15. A simplified version of it is the MiniKlexikon, which features over 1,500 articles aimed at beginning readers. Finally, we also scrape the complete archives of Fluter, a magazine aimed at young adults published by the Federal Agency for Civic Education, which contains a large body of non-fiction. All resources are CC-licensed. Table 1 shows the raw token numbers for all corpora. In sum, we get 16.5M lexical tokens. Before further analysis, we extensively clean and normalize our data (details in Appendix C). We share our dataset on huggingface.
5 Construction distribution analysis
As there are no findings for the distribution of utterance-level construction in German, we conduct our own analysis using spacy. We first split larger paragraphs into individual sentences with the included senter and then annotate these with POS and dependency information. This information serves as the base of our construction annotation procedure. We devise standard construction categories in line with comparable efforts for English Cameron-Faulkner et al. (2003); Cameron-Faulkner and Noble (2013); Bunzeck and Diessel (2024), and use the spacy-annotated data to assign one of the following categories to each utterance:
  • FRA – utterances that do not contain a verb
  • QWH – wh-question (introduced by interrogative pronouns)
  • QYN – yes/no-question (introduced by verbs/auxiliaries)
  • COP – subject-predicate utterance where the predicate is a copula verb (a form of sein or werden)
  • IMP – utterances introduced by verbs in imperative mood
  • SPI – standard subject-predicate utterance (intransitive verb with no direct/accusative object)
  • SPT – standard subject-predicate utterance (transitive verb with direct/accusative object)
  • COM – utterances with two or more lexical verbs
This holistic taxonomy is applicable to every utterance across our subcorpora. Against a manually annotated sample of 1,000 sentences, which we balanced for individual subcorpora, our classifier reaches an accuracy of approx. 95%.
Figure 1 visualizes the results of this corpus annotation process, exact proportions are reproduced in Appendix D. Generally, our results confirm earlier findings (Cameron-Faulkner et al., 2003; Cameron-Faulkner and Hickey, 2011; Cameron-Faulkner and Noble, 2013; Bunzeck and Diessel, 2024): Just like English CDS, German CDS features more questions than any other corpus, abounds with fragments, and contains comparatively few complex utterances. The Project Gutenberg data, on the other hand, is characterized by over 60% complex sentences. Interestingly, the construction distribution forms a continuum across our subcorpora. The MiniKlexikon, for example, contains considerably less complex sentences than the other written genres, but over half of its utterances are (in)transitive, canonical SV-sentences. This shows that even these particular sub-genres of child-directed linguistic input feature highly varied and specific constructional profiles that differ from each other.
6 Training data composition
We compose three different corpora of 5M words: i) one corpus maximally resembling the construction composition of child-directed speech (cds), ii) one corpus containing a drastically higher amount of complex sentences, mirroring the distribution in the Project Gutenberg data (pjg), and iii) a corpus that is averaged between these two (mix). The relative distributions of construction types can be found in Table 2.
Construction cds mix pjg
FRA 25% 16.5% 8%
QWH 9% 5.5% 2%
QYN 21% 12.5% 4%
COP 8% 6.5% 5%
IMP 5% 3.5% 2%
SPI 10% 9% 8%
SPT 12% 11% 10%
COM 10% 35.5% 61%
Table 2: Construction proportions of our training sets
Crucially, we sample the individual utterances for our training sets from all subcorpora in our German BabyLM dataset (cf. Table 1). By doing so, we approximate a similar (if not completely equal) mixture of sources and, therefore also a similar mixture of registers, semantic content, etc. This enables us to isolate the effect of construction distributions in our model’s training data, without any interference from the possible differences between the subcorpora.
Character Subword
cds mix pjg cds mix pjg
Word-level Lexical decision 97.4% 97.6% 97.4% 84.6% 81.9% 80.8%
Surprisal 99.8% 99.8% 99.9% 91.5% 90.3% 90.1%
AntiSurprisal 99.3% 98.9% 99.7% 76.5% 75.4% 75.4%
Syntax Simple Agreement 90% 90% 95.7% 80% 84.3% 92.1%
Across a Prepositional Phrase 61.5% 65.5% 61.8% 74.8% 73.5% 75.5%
Across a Subject Relative Clause 67.1% 66% 62.4% 78.4% 73.7% 97.9%
Short Verb Phrase Coordination 69.8% 68.8% 67.9% 82.6% 93.5% 99.5%
Long Verb Phrase Coordination 53.6% 60.6% 63% 60.6% 78.8% 78%
Across Object Relative Clause 58.6% 54.2% 53% 64% 66.7% 81.6%
Within Object Relative Clause 59.8% 56.4% 72.5% 55.8% 55.7% 49.9%
Semantics XCOMPS 51.5% 49.1% 49.1% 51.4% 52% 52.3%
Table 3: Final evaluation results (accuracies) for all benchmarks
7 Model training and evaluation
We train small Llama models Touvron et al. (2023) with transformers Wolf et al. (2020). To account for the effect of subword tokenization, we compare character-level (3.7M parameters) and subword models (7.7M parameters) for the three datasets. We train all models for one epoch (loss curves and hyperparameters are in Appendix E) and share them on the huggingface model hub ([LINK REMOVED]). To test the effect of different random initializations and our sampling strategy, we reproduce pre-training for the cds models (see Appendix F).
In line with current best practices to linguistic probing, we use minimal pair probing datasets to evaluate our LMs’ linguistic knowledge in German. The datasets always consist of a correct/grammatical and a matched incorrect/ungrammatical string. We use minicons Misra (2022) to score the sentences. We evaluate 19 model checkpoints (10 for the first 10% of training, 9 for the remaining 90%).
Word-level probing
To gauge lexical knowledge, we adapt the experimental setup from Bunzeck et al. (2025): We use wuggy Keuleers and Brysbaert (2010) to generate 1,000 nonce words (e.g. promsen) from existing words (e.g. bremsen) and then evaluate how surprised the models are by i) the words with the context of a prepended white space (lexical decision, Le Godais et al., 2017), ii) the words in a plausible context sequence (surprisal, Hale, 2001), and iii) the words randomly inserted into implausible contexts (antisurprisal, Shafiabadi and Wisniewski, 2025). If the model is less surprised by the existing word, we count this as a correct choice in our paradigm. We calculate accuracies over the whole dataset.
Syntactic probing
For syntactic probing, we use the CLAMS dataset Mueller et al. (2020), which contains syntactic minimal pairs (grammatical/ungrammatical) for German (e.g. Die Autoren lachen/*lacht.). The included seven phenomena all revolve around subject-verb agreement in different contexts (across PPs, relative clauses, with coordination, etc.), resulting in different degrees of difficulty. We score the sentences for their likelihood. We calculate accuracies for correctly rated pairs (grammatical sentence more likely) over the whole dataset.
Semantic probing
To evaluate our models’ semantic knowledge, we use the XCOMPS dataset He et al. (2025). It contains conceptual minimal pairs (e.g. Garnele hat einen Kopf./*Ein Bikini hat einen Kopf.) [^1] We sample 1,000 MPs with randomized replacement, as the other conditions contain implausible/wrong minimal pairs. Furthermore, the quality of translation is not optimal, as exemplified by the missing determiner in front of Garnele. that test whether LMs have acquired knowledge about conceptual properties of real-world entities. Again, we score the sentences for their likelihood and calculate accuracies for correctly rated pairs over the whole dataset.
(a) Trajectories for character models
(b) Trajectories for subword models
Figure 2: Learning trajectories for all minimal pair benchmarks
8 Results
8.1 MP probing
Table 3 shows model-wise accuracies for all minimal pair sets after training for one epoch. For the word-level evaluations, accuracy scores are generally high. Across all tasks, the character models perform with almost perfect accuracy. No effect of the constructional composition of the training data is identifiable here. For the subword models, this is not true. Here, the model trained on more questions/fragments and less complex utterances (cds) outperforms the model that approximates written language on the construction level (pjg). The improvements range from 1% for anti-surprisal to 2-3% on lexical decision.
For the syntactic tests, the picture is more nuanced. Generally speaking, all our models learn to distinguish grammatical and ungrammatical sentences with agreement phenomena. The best scores are achieved on more simplistic phenomena like simple agreement or coordination with short verb phrases. Agreement phenomena that involve longer dependencies and distracting nouns, e.g. within and across relative clauses, are the hardest to learn. For the character models, the cds model outperforms the others on three out of seven tests, including both “across subj./obj. relative clause” conditions. For three other tests, the pjg model wins out, whereas the mix model achieves the highest scores on only one test (agreement across prepositional phrases). The subword models show a somewhat different picture. Here, the pjg model wins in five categories, whereas cds and mix achieve best scores in one each.
The evaluation scores for the XCOMPS benchmark reveal that our small models do not reliably learn the conceptual knowledge underlying the included minimal pairs. Scores revolve around the chance baseline, with subword models performing slightly better than character models for 2/3 data mixtures. However, these scores are also not much worse than those reported in the XCOMPS paper (e.g. 56–69% for Llama 3.1), so more research is needed to establish ways of robust and appropriate probing of semantic knowledge in small LMs.
8.2 Learning trajectories
Figure 2 shows the learning trajectories of our models across one training epoch. In line with best practices in ML Viering and Loog (2023), we log-scale the x-axis in our plots. This allows us to also trace early learning in more detail.
For our character models, word-level learning happens rapidly in an s-shaped curve. No differences are visible between the datasets, performance improvements align almost perfectly. For the subword models, the learning processes are not as nicely monotonically improving. Rather, the learning trajectories show a dip early in training, which then later on recovers to fairly good accuracy scores. Interestingly, despite differences in final scores, the improvements across models trained on quite different datasets still align with regard to turning and takeoff points.
This pattern is also confirmed by the learning trajectories for the syntactic phenomena. While the pjg models trained on more complex utterances frequently reach the highest final scores, it is remarkable to see how the improvements for all models seem to happen in parallel. The global shape of the trajectory is the same for all syntactic tests, regardless of the construction distribution. For example, the learning curve for simple agreement is steeper for the pjg models once learning has started, but take-off points are neatly aligned. These take-off points are pushed back by the individual paradigms’ complexities – simple agreement and short VP coordination begin to improve earlier than MPs containing RCs. Finally, it is interesting to note that for the character models, word-level learning consistently stabilizes before syntactic learning, whereas both processes seem to happen concurrently in subword models (mirroring findings for English reported by Bunzeck and Zarrieß, 2025).
9 Discussion and conclusion
This paper set out to investigate whether the constructional profile of CDS, which is shaped in a way to support the acquisition of functional language competence, is actually conducive to formal language learning, or whether its relative lack of complex sentences and canonical SV(X) utterances makes it so impoverished that meaningful formal learning does not happen. The results of our utterance-level corpus analysis for German align with earlier findings on CDS and book language for English Cameron-Faulkner et al. (2003); Cameron-Faulkner and Noble (2013); Bunzeck and Diessel (2024) and Irish Cameron-Faulkner and Hickey (2011), adding to the growing evidence that this linguistic distribution is fairly universal, at least in WEIRD societies Henrich (2024).
The results of our language modeling experiments are much more surprising. The constructional profile of training data is not overly important for training small, developmentally plausible LMs from scratch. Rather, starting/turning points of the resulting learning trajectories are mostly determined by the respective amount of training steps. Despite models trained with more complex input resulting in slightly better performance, they do not begin to learn earlier. Global learning trajectories are extremely similar, only the local magnitude differs between different constructional setups. For word-level learning processes such as lexical decision or (anti)surprisal tests, data with more fragments and questions even seems to be rather conducive, whereas models trained on more complex utterances do not learn this level of knowledge equally well.
What does this now mean for theories of language acquisition? On a formal level, there seems to be no disadvantage for models trained on allegedly less “complex” or somewhat impoverished data. Despite more complex data leading to slightly better benchmark scores, the learning trajectories remain largely unaffected. What really shapes the learning process is the amount of input, not its formal complexity (similar to findings for children by Huttenlocher et al., 1991; Rowe, 2012). An increase in appropriate construction types for child-rearing and functional language learning (like questions, imperatives, or fragments) does not hinder formal learning (if only reduce its magnitude slightly). Conversely, it even enables word-level learning to converge to a better end state. This also aligns with a broader trend found in language acquisition studies — the complexity and quality of input can indeed predict later language skills Noble et al. (2020); Alroqi et al. (2023), but the ground level is always extremely high already: being a competent user of the language itself. Furthermore, quality varies with many more extralinguistic factors like the number of siblings Laing and Bergelson (2024) or cultural factors Bergelson et al. (2023); Bunce et al. (2024).
Finally, our findings also add to the growing body of research on BabyLMs Warstadt et al. (2023); Hu et al. (2024). Similarly to English models, our German BabyLMs only need little data – the cds dataset contains approx. 820,000 sentences, and given the estimation by Cameron-Faulkner et al. (2003) that children hear around 7,000 utterances per day, our data approximates the number of utterances heard over only 120 days – to learn a fair amount of syntax and almost impeccable lexical knowledge, with trajectories mirroring those of English models Bunzeck and Zarrieß (2025). We hope that our dataset enables other scholars to carry out experiments with developmentally plausible LMs beyond the dominating English LMs, and that our data sources provide inspiration to those compiling BabyLM corpora for their own languages of interest.
Appendix A Limitations and ethical considerations
A.1 Limitations
Our study is limited by data availability. Creating a full-fledged 100M-token BabyLM dataset is currently out of question, as neither CHILDES nor other sources contain even remotely enough data for languages other than English. Principally, synthetic corpora like the TinyStories dataset Eldan and Li (2023), which contains children’s stories generated by GPT-3 or TinyDialogues by Feng et al. (2024) would provide an unlimited source of training data. However, our inspection of their generated dialogues yielded that they drastically underestimate the high numbers of grammatical fragments, questions and short SV(X)-utterances in real-world data. Similarly, there are little to no evaluation sets for German beyond those that we included/creates ourselves, especially on the syntactic level.
Moreover, actual developmental plausibility also hinges on the inclusion of other modalities. For audio data, there are few CHILDES subcorpora and other corpora that contain phonetic information Lavechin et al. (2023), but larger models need to be trained on more data, e.g. audiobooks Lavechin et al. (2025). A middle ground is training on textual phonetic transcriptions generated from raw text, e.g. for the BabyLM data Goriely et al. (2024); Bunzeck et al. (2025). More recently, also video recordings from infant-mounted cameras have been used to train on combined visual and auditory input modalities Wang et al. (2023); Vong et al. (2024); Long et al. (2024). The inclusion of such data could help to disentangle learning processes further.
A.2 Ethical considerations
Given the nature of this work, there are no specific ethical concerns to address. However, we want to emphasize that BabyLMs are not actual babies, but rather abstractions, or models in the original scientific sense, of the distributional, frequency-driven aspects of their learning capacity. All claims regarding their implications for language development in the real world should be understood in this context, which we also attempted to explicate by distinguishing functional and formal aspects of learning.
Appendix B Excluded corpora
Several corpora that are — in principal — available for German were excluded from our analysis. The Folk corpus Reineke et al. (2023) and the Simple German corpus Jach and Dietz (2024) are not available under any open licenses, while the data in other German reference corpora Kupietz et al. (2010) are not available in their entirety but can only be queried through web interfaces. Finally, Homebank features day-long audio recordings of children and their surroundings/inputs VanDam et al. (2016), but without any written transcriptions.
Appendix C Data cleaning
In line with best practices in language modeling, we extensively clean and normalize our data. Our cleaning script is available at [link removed for anonymization].
All subcorpora
We replaced all local variants of single/double quotation marks with either ’ ’ or " ". We further reduced multiple superfluous whitespace and newlines to singular whitespaces.
Talkbank data
For the data sourced from talkbank (i.e. the CHILDES corpora and CallHome), we remove all mark-up and additional info on false starts, hesitations, implicit completions or other explanations. Furthermore, we also remove all empty utterances and those containing xxx or yyy, placeholder symbols for personally identifiable information.
Project Gutenberg
For the Project Gutenberg data, we excluded all lines with more than 6 consecutive whitespaces, as these always turned out to be title pages, index pages, etc., which contain no useful language data. Additionally, we removed all textual data in square brackets, which almost always corresponded to pointers to pictures which are not found in text-only version, or additional explanations by the volunteers who digitized the respective books.
OpenSubtitles
For the OpenSubtitles data, we removed all text in parentheses, which corresponds to speaker information. Also, we removed sentence-initial dashes (-) which were sometimes added. We also amended OCR errors (like mangled uppercase I and lowercase l) as far as possible.
Fluter
For the data sourced from the Fluter magazine, we removed all lines containing additional metatextual data, like author info and image credits, before pre-training.
Appendix D Exact construction proportions
Table 4 shows the exact construction proportions for all of our subcorpora. This data underlies the visualization in Figure 1.
Construction Proj. Gut. Dreamb. Fluter News Wikib. Klex. Mini-Klex. OpenSub. CallHome Child speech CDS
FRA 7.8% 6.3% 6.2% 4.0% 11.6% 6.3% 2.5% 24.1% 37.0% 55.1% 24.5%
QWH 1.9% 0.3% 2.6% 1.4% 0.5% 2.9% <0.1% 7.3% 2.1% 3.5% 8.8%
QYN 3.7% 0.7% 2.8% 1.6% 0.5% 0.4% <0.1% 10.9% 6.9% 4.7% 20.7%
COP 4.6% 7.1% 7.7% 7.4% 10.9% 13.2% 21.4% 9.7% 10.7% 5.7% 8.1%
IMP 1.5% 0.1% 0.2% 0.1% 0.3% <0.1% <0.1% 4.6% 0.4% 2.0% 4.5%
SPI 7.5% 9.2% 9.7% 13.7% 9.5% 13.9% 19.9% 9.9% 8.8% 11.5% 10.1%
SPT 10.5% 14.5% 18.7% 25.7% 24.1% 28.1% 37.2% 18.0% 14.1% 11.9% 12.3%
COM 62.5% 61.8% 52.2% 46.1% 42.7% 35.2% 18.9% 15.4% 20.0% 5.7% 11.0%
Table 4: Exact proportions of constructions for all subcorpora
Appendix E Model hyperparameters and training details
Our models share a hidden/intermediate/embedding size of 256, 8 hidden layers and attentions heads, and a context length of 128. For the character models, the vocabulary consists of all printable ASCII characters and characters used in written German (üäöß and their uppercase variants), amounting to a vocab. size of 110 and 3,730,688 parameters. For the subword models, we train a BPE tokenizer Gage (1994) with a vocab. size of 8,000 and add two special tokens (BOS, EOS/PAD), resulting in 8,002 vocab. tokens and 7,771,392 parameters. Model training takes approx. 2h on a MacBook Pro with an Apple M2 Pro CPU/GPU.
We reproduce the training and test loss curves for our models in Figure 3. For the test loss, we evaluated perplexity on a held-out, randomly sampled portion of each individual training corpus. We find no principal differences in loss development, although the character models and models trained on the cds data seem to converge the fastest. As the similar curves for train and test loss indicate, all models succeed in optimizing for their next-token prediction goal.
Figure 3: Loss curves for our self-trained character and subword models
Appendix F Repeated training runs
A common criticism in the BabyLM paradigm is the purported effect of training noise on model performance, which is hard to disentangle from real training data effects. While training and evaluating multiple random seeds for all our models would be too costly, we repeated two additional training runs for the character-level cds model with different random initializations (learning trajectories in Figure 4(a)) and two additional training runs where we re-sampled the cds dataset from our whole corpus with the exact same construction composition, but different content (learning curves in Figure 4(b)). In both cases, the learning trajectories do not differ tremendously. For the word-level phenomena (LexDec, Surprisal, AntiSurprisal), the curves overlap almost perfectly. For the syntax phenomena, we can see some variation and oscillation in the curves, but the trajectories still remain extremely similar (and do not differ in their steepness, the main effect that we see in Figure 2 between the datasets with different construction compositions).
(a) Trajectories for different random initializations
(b) Trajectories for different samples of cds data
Figure 4: Learning trajectories for our comparison models